OpenStudy (anonymous):

What is the inverse of y = (1/2)x + 3?

OpenStudy (anonymous):

switch x and y x = (1/2)y + 3 y = 2(x - 3)

OpenStudy (lukecrayonz):

Solve for x, then replace y with x :)

OpenStudy (lukecrayonz):

Or replace y with x in the beginning, either way.

OpenStudy (anonymous):

Thx